40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4301     Entry time: Tue Feb 15 11:57:06 2011
Author: steve, valera 
Type: Configuration 
Category: PSL 
Subject: PMC swap 

 We swapped the PMC s/n 2677 for s/n lho006.

The table below summarizes the power levels before and after the PMC swap.

  old new
Ptrans 1.32 W 1.42 W
Transmission 85 % 91.5 %
Refl PDDC locked/unlocked 5.0 %  4.3 %
Loss 7-8 % 2-3 %
Leakage out of the back 10 mW 0.3 mW

 

- The power into the PMC (1.67 W) was measured with Scietech bolometer before the first steering PMC mirror. The leakage through the steering mirrors was measured with Ophir power meter to be 12+8 mW. There is also a lens between the mirrors which was not measured. 

- The power through the PMC was measured after the doubler pick off (105 mW), steering mirror (4 mW), and lens (not measured).

- The estimated reflection from four lens surfaces is 1-2% hence 1% uncertainty in the losses in the table.

- The beams into the PMC and on REFL PD were realigned. The beams downstream of the PMC are blocked as we did not realigned the PMC and doubler paths.

- The trans PD ND filters were removed. The VDC=1.28 V now.

- The NPRO current is 2.102 A

 

Atm 1 old

Atm2  new

Attachment 1: P1070421.JPG  2.965 MB  | Hide | Hide all
P1070421.JPG
Attachment 2: P1070423.JPG  2.969 MB  | Hide | Hide all
P1070423.JPG
ELOG V3.1.3-